# The Labyrinth of Links: Navigating the Associative Maze of Multi-modal LLMs

### v1

@@ 5.3.1 Single-step Association
Table 1 shows the mean result in adjective and verb concepts, i.e., attribute and affordance on the OCL Li et al. (2023) and verb on the Pangea Li et al. (2024b). We focus on the metric of “success ratio” in this part, and the detailed result on each item can be found in supplementary Table 5,  6,  7.
From the result, we find that Gemini-1.5-Flash achieves the highest performance across various concepts in all settings, but remains a certain gap from humans, i.e., success ratio compared to $1$. It deserves noted that human single-step associations are nearly unlikely errors, as correct memory entails the rule. Moreover, our comprehensive data refinement ensures the data quality that prevents any errors induced by annotations confusion. Besides, the association has certain improvements across various concepts, i.e., the proposed memory strategy is higher than NoM. Simultaneously, the NLM attained the strongest performance, we speculate that due to the current MLLM being trained on a large amount of internet visual language data, lacking structured data training.
@@ 5.4 Analysis of Results
We now analyze the failure cases in the association process. We divided them into two types according to the stage in our benchmark, i.e., Error deduction, and Error Association. In the first type, the MLLM produces the error perception about the common concepts in two samples. This then causes the error in the association step (Left of Figure 5). Conversely, in the second type, MLLM has the correct memory and makes errors due to the limited perception capability (Right of Figure 5). More interesting, these failures are consistent with human tests, we may derive error information from examples and further induce incorrect judgments.
@@
We demonstrate that there is a significant gap between current open-source, and closed-source MLLMs and human intelligence. We speculate that the underperformance arises from two main factors: a lack of multi-image instruction tuning and limitations in contextual understanding. The first statement is also observed from existing work Song et al. (2024); Wang et al. (2024c), that the current MLLM has a weak ability in multi-image understanding since the lack of its instruction data. In addition, recent research Wang et al. (2024a) has demonstrated MLLMs’ poor performance in handling long contexts, i.e., memory. We also visualize the attention map of open-source MLLM Qwen-VL on OCL attribute concept to support this observation, as Figure 6, we can easily find that response has predominate attention at the position close to response instead of the part for decision-making in both StructM and NLM.
@@ A.1 Implementation of Data Refinement
In our experiment, we comprehensively develop three steps to consolidate the data quality in our tasks, including Image resolution filter, MLLM verification, and Human expert evaluation.
@@
The first two steps ensure the correctness of our benchmark, which reduces the image with low quality or confusing annotation. We continue to develop an online testing interface for human testers, as shown in Figure 7, which further filters the images with potential confusion or ethical concerns from the human perspective. We have strictly followed the ethical review, which is described in section F of the supplementary.
@@ A.2 Implementation of Memory Base
Appendix B Preliminary investigation on Concept Perception
@@
We carry out experiments on three open-source MLLMs. QWen-VL Bai et al. (2023b) uses QWen-7B Bai et al. (2023a) as the initialize of LLM and OpenCLIP’s ViT-bigG Cherti et al. (2023) as the visual encoder. LLaVA-NeXT-7B Liu et al. (2024) is constructed on the LLM of Mistral-7B Jiang et al. (2023) and vision encoder of CLIP’s ViT-L/14 Radford et al. (2021), while LLaVA-NeXT-13B uses the same vision encoder but is constructed on LLM of Vicuna-13B Zheng et al. (2024).
@@
Appendix C Detailed Results
C.1 Detailed Single-step Association Result
@@
NLM $65.26$ 93.06 $\underline{96.27}$ 97.83 $52.61$ $\underline{96.41}$ $92.91$ $98.65$ $86.63$
ChainM $63.90$ $91.59$ $92.55$ $93.48$ $55.15$ 97.01 $93.70$ $1.00$ $73.55$
@@
StructM $68.8$ $91.8$ $93.7$ $\underline{97.8}$ 58.4 $92.7$ 95.2 100.0 $\underline{87.30}$
NLM 74.9 $\underline{92.9}$ 96.8 $95.5$ $\underline{57.5}$ $94.8$ $93.9$ 100.0 88.39
ChainM $\underline{70.3}$ $92.6$ $87.1$ $90.9$ $56.4$ $79.9$ $\underline{94.4}$ $98.6$ $83.78$
Deduction Success Ratio LLaVA-Onevision - $\underline{71.73}$ $39.40$ $38.51$ $\underline{69.57}$ $17.43$ $\underline{55.69}$ $\underline{90.55}$ $12.16$ $49.38$
QWen2-VL - $63.10$ 69.75 $51.55$ $63.04$ $\underline{39.62}$ 59.88 $59.84$ $58.11$ $58.11$
mPLUG-Owl3 - $66.76$ $44.12$ $\underline{61.49}$ 84.78 $32.49$ $47.31$ $71.65$ $81.08$ $\underline{61.21}$
Gemini-1.5-Flash - 74.44 $\underline{62.18}$ 79.75 84.78 52.10 $6.70$ 91.60 $75.00$ 65.82
@@
StructM 97.02 $70.72$ $\underline{91.77}$ $\underline{73.99}$ $\underline{98.19}$ $58.33$ 99.04 86.55 $\underline{84.45}$
NLM 97.02 83.64 92.90 74.79 98.20 $55.17$ $\underline{98.06}$ $83.44$ 85.40
ChainM $\underline{94.64}$ $76.50$ $86.29$ $66.40$ $95.58$ $46.67$ $95.19$ $78.92$ $80.02$
@@
mPLUG-Owl3 - 82.94 $0.84$ $\underline{12.12}$ $\underline{2.89}$ $63.91$ $\underline{12.90}$ 84.91 $\underline{5.87}$ 33.30
@@
Gemini-1.5-Flash NoM $85.95$ $88.96$ $\underline{89.23}$ $71.88$ $94.16$ $81.94$ $92.05$ $69.47$ $84.21$
StructM $86.49$ $89.81$ $85.71$ $\underline{77.97}$ 99.60 $\underline{93.20}$ $\underline{97.72}$ $74.33$ $\underline{88.10}$
NLM 93.07 $\underline{92.36}$ 92.28 $75.41$ $98.99$ 93.78 $97.28$ $73.48$ 89.58
ChainM $86.46$ 93.54 $88.34$ $74.63$ 99.60 $86.23$ 98.41 $73.42$ $87.58$
Deduction Success Ratio LLaVA-Onevision - $58.87$ $\underline{48.97}$ $73.35$ $34.78$ $\underline{96.85}$ $3.96$ $3.60$ $\underline{52.48}$ $46.61$
QWen2-VL - $\underline{60.90}$ 69.00 $67.41$ $\underline{62.32}$ $89.78$ $7.64$ $3.60$ $39.26$ $\underline{49.99}$
mPLUG-Owl3 - $33.99$ $4.94$ $\underline{74.52}$ $44.93$ $96.74$ $\underline{7.72}$ $\underline{5.63}$ 74.62 $42.89$
Gemini-1.5-Flash - 67.21 $18.00$ 86.99 79.69 97.79 19.96 59.68 $16.22$ 55.69
@@
C.2 Detailed Synchronous Association Result
@@
C.3 Detailed Asynchronous Association Result
@@
C.4 Detailed Ablation Study
@@
Table 14 shows the detailed results of the ablation study of example sizes. We observed that while the average max and mean step slightly differ between different example sizes, there are notable gaps within specific paired categories. For instance, in the case of LLaVA-OneVision with paired attribute “furry-metal”, StructM-3 outperforms StructM-5 by approximately the mean-step of $1.95$. Furthermore, while an average mean-step of $3$ outperforms in most cases, the example size of $1$ and $5$ also leads in certain cases. For instance, with QWen2-VL on “painted-rusty” affordance and LLaVA-OneVision on “fresh-cooked” attribute, the example size of 5 outperforms the others.
Appendix D Comparison with More Concepts
@@
Table 15: The Max $\mid$ Mean Step on paired action concept association from HMDB dataset in asynchronous association setting. We implemented at open-source MLLM QWen-VL, close-sourced MLLM Gemini-Pro-Vision and GPT4-V, as well as human experts. The best and second results are shown in bold and underline, respectively.
@@
Appendix E Comparison with unfiltered Raw Data

### v2

@@ 5.3.1 Single-step Association
Table 1 shows the mean result in adjective and verb concepts, i.e., attribute and affordance on the OCL Li et al. (2023) and verb on the Pangea Li et al. (2024c). We focus on the metric of “success ratio” in this part, and the detailed result on each item can be found in supplementary Table 5,  6,  7.
Type Concept Memory Models (Success Ratio)
LLaVa-OneVision QWen2-VL mPLUG-Owl3 Gemini-1.5-Flash GPT-4o Avg.
Association Success Ratio Attribute (Adjective) NoM $75.52$ $76.36$ $54.05$ $75.40$ $84.49$ $73.16$
StructM $78.41$ $77.74$ $72.67$ $87.30$ $88.83$ $80.39$
NLM $80.40$ $86.63$ $73.33$ $88.39$ 89.87 $83.72$
ChainM $77.73$ $73.55$ $70.70$ $83.78$ $78.86$ $76.92$
Affordance (Adjective) NoM $70.61$ $75.10$ $53.80$ $77.78$ $84.93$ $72.44$
StructM $73.39$ $75.39$ $73.02$ $84.45$ $85.36$ $61.25$
NLM $76.13$ $79.37$ $67.79$ $\underline{85.40}$ 86.76 $79.10$
ChainM $74.38$ $82.67$ $68.46$ $80.02$ $81.54$ $77.40$
Action (Verb) NoM $75.74$ $78.43$ $57.94$ $84.21$ $86.97$ $76.66$
StructM $75.44$ $82.10$ $73.92$ $88.10$ $\underline{88.72}$ $81.66$
NLM $78.66$ $88.01$ $70.04$ 89.58 $86.13$ $82.48$
ChainM $76.92$ $85.59$ $69.50$ $87.58$ $85.90$ $81.10$
Deduction Success Ratio Attribute (Adjective) $49.38$ $58.11$ $61.21$ $65.82$ $78.14$ $62.53$
Affordance (Adjective) $21.07$ $15.98$ $33.30$ $33.08$ $27.30$ $26.15$
Action (Verb) $46.61$ $49.99$ $42.89$ $55.69$ $57.28$ $50.60$
The result shows that GPT4-o achieves the highest performance in adjectives and Gemini-1.5-Flash makes advances in verbs, but those all remain a certain gap from humans, i.e., success ratio compared to $1$. It deserves noted that humans are unlikely errors as association links are given in context. Moreover, our comprehensive data refinement ensures the data quality that prevents any errors induced by annotations confusion. Besides, the association has certain improvements across various concepts, i.e., the proposed memory strategy is higher than NoM. Simultaneously, the NLM attained the strongest performance, we speculate that due to the current MLLM being trained on a large amount of internet visual language data, lacking structured data training.
@@ 5.4 Analysis of Results
In the following, we first discuss some failure cases and then deeply analyze the potential reasons why MLLMs are inferior to humans via attention and underlying capabilities  [^5] Due to limited space, the detailed analysis refers to Section C in the supplementary..
@@
We divided the errors into two types according to the stage in our benchmark, i.e., Error deduction, and Error Association. For error deduction, the MLLM predicts the error links for associated objects. This then causes the error in the association step (Left of Figure 5). Conversely, for error association, MLLM has the correct memory and makes errors due to the limited perception capability (Right of Figure 5). More interestingly, these failures are consistent with humans, we may derive error information and further induce incorrect judgments.
@@
We demonstrate that there is a significant gap between MLLMs and humans. We speculate their two main factors: a lack of multi-image instruction tuning and limitations in contextual understanding. The first statement is also observed from existing work Song et al. (2024); Wang et al. (2024c), that the current MLLM has a weak ability in multi-image understanding since the lack of its instruction data. In addition, recent research Wang et al. (2024a) has demonstrated MLLMs’ poor performance in handling long contexts, i.e., memory. We also visualize the attention map of open-source MLLM Qwen-VL on OCL attribute concept to support this observation, as Figure 6, we can easily find that response has predominate attention at the position close to response instead of the part for decision-making in both StructM and NLM.
@@ 7 Conclusion
Acknowledgments
This work is supported in part by the National Natural Science Foundation of China under Grant No.62306175, CCF-Tencent Rhino-Bird Open Research Fund.
@@ A.1 Implementation of Data Refinement
Figure 7: Annotation-free construction method. The left transfers raw supervised data for the association dataset. The right is the expansion of three strict data quality control steps, including image resolution filter, MLLM verification, and human expert evaluation.
In our experiment, we comprehensively develop three steps to consolidate the data quality in our tasks, including Image resolution filter, MLLM verification, and Human expert evaluation, as seen in Figure 7.
@@
The first two steps ensure the correctness of our benchmark, which reduces the image with low quality or confusing annotation. We continue to develop an online testing interface for human testers, as shown in Figure 8, which further filters the images with potential confusion or ethical concerns from the human perspective. We have strictly followed the ethical review, which is described in section H of the supplementary.
@@ A.2 Implementation of Memory Base
A.3 Implementation of Evaluation Setting
Transferring from the single-step to synchronous and asynchronous steps, we evaluate the association capability in a dynamic sequential environment. This involves iteratively processing the association step and deduction step and exiting when the association makes an error. In this setting, we evaluate the max / mean step, which is the maximum step and average step across multiple round tests, respectively. The related evaluation setting can refer to Algorithm 1.
To save memory during synchronous and asynchronous steps, we developed three distinct strategies for transferring experiences in the deduction step to update association memory. These strategies encompass memory storage, deep memory retention, and forgetting mechanisms.
Algorithm 1 Synchronous/Asynchronous Association Evaluation
1:procedure Main
2:     for all attr $\in$ options do
3:         for epoch $\leftarrow$ 1 to $N$ do
4:              Reset_Memory()
5:              for i $\leftarrow$ 1 to max_len do
6:                  if Asynchronous and i mod 5 = 4 then
7:                       Switch_Attribute() $\triangleright$ Change to the other attribute in Asynchronous.
8:                  end if
9:                  query_img $\leftarrow$ previous_correct_img
10:                  correct_img, false_img $\leftarrow$ Get_Next_Images(attr)
11:                  correct_answer $\leftarrow$ Random_Order() $\triangleright$ Randomly set Img1 or Img2 as correct image
12:                  response $\leftarrow$ Query_MLLM(memory, query_img, correct_img, false_img)
13:                  if response $\neq$ correct_answer then
14:                       Record_Step_Length(i)
15:                       break
16:                  end if
17:                  if mode $\neq$ “nomem” then
18:                       Deduction_Step(query_img, correct_img)
19:                  end if
20:              end for
21:         end for
22:         Calculate_Statistics(steps) $\triangleright$ Calculate $Max|Mean$ step for all epochs.
23:     end for
24:end procedure
Appendix B Deeply Comparison with Existing Benchmark
Our paper is the first to conduct a comprehensive investigation into sequential concept association in MLLMs. - As summarized in a recent benchmark survey Li et al. (2024b), MLLMs’ benchmark mainly concentrates on the design of more complex tasks and evaluating nuanced correlations between input samples. - More deep comparison, our benchmark also needs general perception capability as LLaVA-Bench Liu et al. (2024b), nuanced features within images as Compbench Kil et al. (2024), and cooperation across different modalities as SpatialRGPT Cheng et al. (2024). Furthermore, our benchmark needs exceptional abilities beyond existing work. First, our benchmark builds on adjectives and verbs that need deeper perception than nouns in MMVP Tong et al. (2024). Second, our benchmark is sequential images beyond the two images in the existing work of MILEBENCH Song et al. (2024). Finally, our benchmark breaks the closed reasoning into open scenarios, which evaluate an inductive and deductive process rather than ground-truth ’yes/no’ in existing work as MARVEL Jiang et al. (2024).
In summary, our benchmark involves the basic capability existing in the previous benchmark, including general perception, nuance features, and cooperation across different modalities. Simultaneously, our benchmark involves exceptional abilities, including deep concept perception, sequential image tasks, and larger solution space.
Figure 9: Comparison with existing work. Existing work can be roughly divided into perception & understanding and cognitive & reasoning. Those all target the design of more complex tasks or nuanced evaluation. Conversely, our work concentrates on evaluating sequential concept association between different objects.
Appendix C Deep analysis of MLLM inferior Humans
We speculate three possible reasons contributing to the significant gap between MLLMs and humans. The related content can be referred to Figure 10.
The benchmark needs more precise locationality.
Compared to existing methods Wu et al. (2024) that rely on language questions with explicit information to retrieve related image content, our benchmark includes a comprehensive memory that stores structured knowledge from prior experiences. The key information relevant to the model’s predictions is often distributed across different locations and represented as relatively local, subtle elements within the context. This highlights the need for more advanced attention in our benchmark.
The association needs more common sense-based reasoning.
The association is an implicit link between the different objects that are not explicitly shown in the query, requiring a deeper common sense understanding. For instance, existing benchmark Song et al. (2024) infer the weather in the image, they directly use the word in the query to retrieve the part in the image. In contrast, our benchmark only specifies the direction of the link, requiring the model to think of the underlying link within the prompted scope.
The association needs short-term memory, which is underexplored by existing work.
Our association benchmark is sequential concept links between different objects, which inherently need the memory mechanism to retain the prior experiences. This aspect of memory remains underexplored in existing benchmarks, with current designs falling significantly behind the sophistication of human cognitive systems. This inspires the future study in memory design.
Figure 10: Deep analysis of MLLM inferior humans. We speculate there exist three possible reasons for this gap, including salience attention, common sense-based reasoning, and human-like memory.
Appendix D Preliminary investigation on Concept Perception
@@
We carry out experiments on three open-source MLLMs. QWen-VL Bai et al. (2023b) uses QWen-7B Bai et al. (2023a) as the initialize of LLM and OpenCLIP’s ViT-bigG Cherti et al. (2023) as the visual encoder. LLaVA-NeXT-7B Liu et al. (2024a) is constructed on the LLM of Mistral-7B Jiang et al. (2023) and vision encoder of CLIP’s ViT-L/14 Radford et al. (2021), while LLaVA-NeXT-13B uses the same vision encoder but is constructed on LLM of Vicuna-13B Zheng et al. (2024).
@@
Appendix E Detailed Results
E.1 Detailed Single-step Association Result
@@
NLM $65.26$ $93.06$ $96.27$ $\underline{97.83}$ $52.61$ $96.41$ $92.91$ $98.65$ $86.63$
ChainM $63.90$ $91.59$ $92.55$ $93.48$ $55.15$ $\underline{97.01}$ $93.70$ $1.00$ $73.55$
@@
StructM $68.80$ $91.80$ $93.70$ $97.80$ 58.40 $92.70$ $95.20$ 100.0 $87.30$
NLM $\underline{74.90}$ $92.90$ $\underline{96.80}$ $95.50$ $\underline{57.50}$ $94.80$ $93.90$ $100.0$ $88.39$
ChainM $70.30$ $92.60$ $87.10$ $90.90$ $56.40$ $79.90$ $94.40$ $98.6$ $83.78$
GPT-4o NoM $58.06$ $89.07$ $96.13$ $90.47$ $52.67$ $95.73$ $95.12$ $98.65$ $84.49$
StructM $74.11$ 98.98 99.39 $97.83$ $44.95$ 99.40 $\underline{96.00}$ $100.0$ $\underline{88.83}$
NLM 78.35 $\underline{96.98}$ $96.20$ $97.83$ $53.77$ $95.83$ 100.0 100.0 89.87
ChainM $72.13$ $90.59$ $81.82$ 100.0 $46.15$ $81.48$ $66.67$ $100.0$ $79.86$
Deduction Success Ratio LLaVA-Onevision - $\underline{71.73}$ $39.40$ $38.51$ $69.57$ $17.43$ $\underline{55.69}$ $\underline{90.55}$ $12.16$ $49.38$
QWen2-VL - $63.10$ $\underline{69.75}$ $51.55$ $63.04$ $39.62$ 59.88 $59.84$ $58.11$ $58.11$
mPLUG-Owl3 - $66.76$ $44.12$ $61.49$ 84.78 $32.49$ $47.31$ $71.65$ $81.08$ $61.21$
Gemini-1.5-Flash - 74.44 $62.18$ $\underline{79.75}$ 84.78 $\underline{52.10}$ $6.70$ 91.60 $75.00$ $\underline{65.82}$
GPT-4o - $59.39$ 86.34 94.19 $73.81$ 90.77 $52.80$ $86.99$ $80.82$ 78.14
@@
StructM $97.02$ $70.72$ $91.77$ $73.99$ $98.19$ $58.33$ 99.04 $\underline{86.55}$ $84.45$
NLM $97.02$ $\underline{83.64}$ $92.90$ $74.79$ $98.20$ $55.17$ $98.06$ $83.44$ $\underline{85.40}$
ChainM $94.64$ $76.50$ $86.29$ $66.40$ $95.58$ $46.67$ $95.19$ $78.92$ $80.02$
GPT-4o NoM $88.54$ $77.58$ $92.47$ $75.64$ $\underline{98.96}$ $60.98$ $95.92$ 89.68 $84.93$
StructM 98.21 87.05 $94.97$ $86.46$ $74.59$ $59.02$ $\underline{98.11}$ $\underline{84.46}$ $85.36$
NLM $\underline{97.65}$ $80.28$ $\underline{96.97}$ 86.76 100.0 $55.56$ $96.00$ $80.82$ $86.76$
ChainM $87.50$ $78.26$ 100.0 $83.33$ $97.99$ $58.06$ $68.86$ $78.28$ $81.54$
@@
mPLUG-Owl3 - 82.94 $0.84$ $\underline{12.12}$ $2.89$ $\underline{63.91}$ $12.90$ 84.91 $\underline{5.87}$ 33.30
@@
Gemini-1.5-Flash NoM $85.95$ $88.96$ $89.23$ $71.88$ $94.16$ $81.94$ $92.05$ $69.47$ $84.21$
StructM $86.49$ $89.81$ $85.71$ $77.97$ 99.60 $\underline{93.20}$ $97.72$ $74.33$ $88.10$
NLM 93.07 $92.36$ $\underline{92.28}$ $75.41$ $98.99$ 93.78 $97.28$ $73.48$ 89.58
ChainM $86.46$ $\underline{93.54}$ $88.34$ $74.63$ 99.60 $86.23$ $\underline{98.41}$ $73.42$ $87.58$
GPT-4o NoM $86.75$ $89.24$ 92.86 $75.56$ $95.34$ $82.56$ $96.95$ $\underline{76.51}$ $86.97$
StructM $81.87$ 96.39 $91.62$ $73.52$ $99.42$ $88.44$ 100.0 78.53 $\underline{88.72}$
NLM $87.37$ $89.41$ $89.12$ 81.97 $99.00$ $80.90$ $90.50$ $70.74$ $86.13$
ChainM $85.00$ $90.40$ $91.00$ $67.65$ $98.99$ $86.29$ $97.99$ $69.85$ $85.90$
Deduction Success Ratio LLaVA-Onevision - $58.87$ $48.97$ $73.35$ $34.78$ $96.85$ $3.96$ $3.60$ $52.48$ $46.61$
QWen2-VL - $\underline{60.90}$ $\underline{69.00}$ $67.41$ $\underline{62.32}$ $89.78$ $7.64$ $3.60$ $39.26$ $49.99$
mPLUG-Owl3 - $33.99$ $4.94$ $74.52$ $44.93$ $96.74$ $7.72$ $5.63$ 74.62 $42.89$
Gemini-1.5-Flash - 67.21 $18.00$ 86.99 79.69 $\underline{97.79}$ $\underline{19.96}$ 59.68 $16.22$ $\underline{55.69}$
GPT-4o - $38.50$ 70.20 $\underline{83.50}$ $58.82$ 100.0 21.83 $\underline{31.66}$ $\underline{53.27}$ 57.28
@@
E.2 Detailed Synchronous Association Result
@@
E.3 Detailed Asynchronous Association Result
@@
E.4 Detailed Ablation Study
@@
Table 14 shows the detailed results of the ablation study of example sizes. We observed that while the average max and mean step slightly differ between different example sizes, there are notable gaps within specific paired categories. For instance, in the case of LLaVA-OneVision with paired attribute “furry-metal”, StructM-3 outperforms StructM-5 by approximately the mean-step of $1.95$. Furthermore, while an average mean step of $3$ outperforms in most cases, the example size of $1$ and $5$ also leads in certain cases. For instance, with QWen2-VL on “painted-rusty” affordance and LLaVA-OneVision on “fresh-cooked” attribute, the example size of 5 outperforms the others.
Appendix F Comparison with More Concepts
@@
Table 15: The Max $\mid$ Mean Step on paired action concept association from HMDB dataset in asynchronous association setting. We implemented at open-source MLLM QWen-VL, close-sourced MLLM Gemini-Pro-Vision, and GPT4-V, as well as human experts. The best and second results are shown in bold and underline, respectively.
@@
Appendix G Comparison with unfiltered Raw Data